### v1

@@ 2.2 Natural Product Taxonomy Classification
We first fine-tuned the model at the Class level. Multiple pre-trained molecular models were used as baselines, replicated, and fine-tuned on this task for comparative analysis. We also selected the widely used non-pretrained representation, Extended-Connectivity Fingerprint(ECFP), as an additional baseline. When splitting the dataset into training, validation, and test sets, we ensured data balance by stratified splits based on the classes. During the comparison, we limited the number of molecules per class in the training set to simulate the classification performance of models under datasets of varying sizes. As shown in Table 1, it can be observed that under different testing dataset sizes, NaFM demonstrates significant performance improvements compared to other baselines, with even more pronounced enhancements when the dataset is smaller. The performance from graph neural network based methods, e.g. MolCLR, PretrainGNN, and D-MPNN express much more fine-tuning data size dependence as their performances drop drastically as we decreased the samples for fine-tuning. GraphMVP has achieved reasonable performance by incorporating implicit 3D information, while Mole-BERT benefits from its prior tokenizer training, which divides atoms of the same type into multiple distinct tokens. This approach increases the number of categories, thereby raising the difficulty of the masked atom modeling task. Overall results suggest that NaFM is the only pretraining model that outperforms rule-based fingerprint ECFP, which is the previous common choice for natural product classification.
@@ 2.4 Natural product bio-activity prediction and virtual screening
The reason for the relatively poor performance of GraphMVP can be attributed to its use of molecular geometry information during pre-training. Given the structural complexity of natural products, the 3D atomic coordinates and bond lengths/angles predicted by RDKit are often inaccurate. This introduction of erroneous geometry information during training can negatively impact the model’s performance, making it less effective at handling the unique complexities of natural product structures.

### v2

@@ 2.2 Natural Product Taxonomy Classification
We first fine-tuned the model at the Class level. Multiple pre-trained molecular models were used as baselines, replicated, and fine-tuned on this task for comparative analysis. We also selected the widely used non-pretrained representation, Extended-Connectivity Fingerprint(ECFP), as an additional baseline. When splitting the dataset into training, validation, and test sets, we ensured data balance by stratified splits based on the classes. During the comparison, we limited the number of molecules per class in the training set to simulate the classification performance of models under datasets of varying sizes. As shown in Table 1, it can be observed that under different testing dataset sizes, NaFM demonstrates significant performance improvements compared to other baselines, with even more pronounced enhancements when the dataset is smaller. Graph neural network-based methods, e.g. MolCLR, PretrainGNN, and D-MPNN, exhibit a strong dependence on fine-tuning data size, with their performance dropping drastically as the number of fine-tuning samples decreases. In contrast, Mole-BERT benefits from its prior tokenizer training, which distinguishes atoms of the same type by assigning them to multiple distinct tokens. Overall results suggest that NaFM is the only pretraining model that outperforms rule-based fingerprint ECFP, which is the previous common choice for natural product classification.
@@ 2.4 Natural product bio-activity prediction and virtual screening
The reason for the relatively poor performance of GraphMVP can be attributed to its use of molecular geometry information during pre-training. As our analysis in Supplementary Section 2.2 indicates, different natural products can exhibit both rigid substructures and highly flexible global conformations, leading to a broader range of possible 3D structures. This inherent flexibility presents a challenge for GraphMVP, as natural products possess a highly diverse 3D distribution, making it difficult for the model to effectively learn and capture their structural variations.
@@ 4.4 Training Details
4.5 Ablation Study
Ablation Study on Model Architecture.
To assess the individual contributions of contrastive learning and masked modeling within the pretraining framework, we performed two additional pretraining experiments by selectively disabling each component. Specifically, we independently ablated the contrastive learning module and the masked modeling module, retraining the model in each case. The results, summarized in Supplementary Table S3, demonstrate the complementary roles of these two components and highlight their respective impact on the overall performance of the pretrained model. It is evident that removing either of the pretraining modules leads to suboptimal results. Notably, the performance gap between the model without contrastive learning and the full NaFM model is the largest among the ablations, underscoring the particularly significant role of contrastive learning in enhancing model performance.
Ablation Study on Contrastive learning
We conducted a comprehensive ablation study to investigate the effectiveness of the NaFM pretraining strategy specifically designed for natural products. Among its core components, contrastive learning plays a critical role, for which we introduced several tailored improvements. As previously discussed, we incorporated scaffold similarity as weighting in the contrastive loss function, and dynamically adjusted the loss weights throughout the pretraining process using a scheduling mechanism. This dynamic adjustment was found to effectively mitigate overfitting in contrastive learning, with the cosine schedule yielding the best results.
To assess the individual contributions of these design choices, we performed a series of ablation experiments, as summarized in Supplementary Table S4. The results demonstrate that removing either the scaffold-based weighting or the loss schedule leads to suboptimal performance. Furthermore, alternative scheduling strategies such as exponential or logarithmic schedules produced slightly inferior results compared to the cosine schedule used in NaFM. These findings highlight that the superior performance of NaFM stems from the synergistic integration of multiple fine-grained enhancements in the contrastive learning component.
Ablation Study on the Impact of Pretraining Dataset.
To isolate the effect of dataset composition and assess the contribution of our methodological innovations independent of data bias, we conducted a dedicated ablation study. Specifically, we examined whether the observed superiority of NaFM could be attributed primarily to the use of a natural product-specific dataset during pretraining. To this end, we selected MolCLR—a representative baseline employing conventional contrastive learning—and pretrained it on the same COCONUT natural product dataset as NaFM. The results presented in Supplementary Table S5 indicate that the newly pretrained MolCLR, despite being trained on the same natural product-specific COCONUT dataset, performs even worse on downstream tasks compared to its original version trained on a general molecular dataset. This suggests that for pretraining methods not specifically tailored to the unique characteristics of natural products, data bias introduced by natural product-specific datasets can actually be detrimental. These findings align with our intuition: without architectural or algorithmic adaptation, simply switching to a domain-specific dataset does not necessarily yield performance gains. Therefore, the superior performance of NaFM cannot be solely attributed to dataset composition; rather, it underscores the necessity of a tailored pretraining strategy to effectively handle the structural complexity and diversity inherent to natural products.